40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14919     Entry time: Tue Oct 1 18:35:12 2019
Author: gautam 
Type: Update 
Category: General 
Subject: Beam centering campaign 
  1. With TRX and TRY maximized using ASS, I centered the Oplev spots on the respective QPDs for the four test masses and the BS. I also centered the spot onto the IPPOS QPD by moving the available steering mirror.
  2. At EX, I tweaked the input pointing of the green beam into the arm by manually twiddling with the PZT mirrors. I was able to get GTRX~0.4.
  3. On the AS table - Koji and I found that there was a steering mirror placed in the AS beam path such that there was no light reaching the AS110 or AS55 PDs. Please - when you are done with your measurement, return the optical configuration to the state it was in before so that the usual locking activity isn't disturbed by a needless few hours troubleshooting electronics.

Once Koji is done with his checkout of the whitening electronics, I will try and lock the PRMI.

ELOG V3.1.3-